Consider a linear regression model where y represents the response variable, xis a quantitative explanatory variable, and d is a dummy variable. The model is estimated as y^  = 15.8 + 4.3x − 3.4d.

a. Interpret the dummy variable coefficient.

  • Intercept shifts down by 3.4 units as d changes from 0 to 1.

  • Slope shifts down by 3.4 units as d changes from 0 to 1.

  • Intercept shifts up by 3.4 units as d changes from 0 to 1.

  • Slope shifts up by 3.4 units as d changes from 0 to 1.

b. Compute y^ for x = 3 and d = 1. (Round your answer to 1 decimal place.)

c. Compute yˆy^  for x = 3 and d = 0. (Round your answer to 1 decimal place.)
Answer #1

a) Intercept shifts down by 3.4 units as d changes from 0 to 1.

Option A is correct.

b) = 15.8 + 4.3*3 - 3.4*1 = 25.3

c) = 15.8 + 4.3*3 - 3.4*0 = 28.7